OpenStudy (anonymous):

Help Me Please Five Question's 6th Grade :D

OpenStudy (anonymous):

Find the least common denominator (LCD) for the fractions. 2/7 and 9/10 A. 1 B. 10 C. 35 D. 70

OpenStudy (anonymous):

2. (1 pt) Choose the missing number to make equivalent fractions. 7/? = 28/40 A. 10 B. 19 C. 28 D. 160

OpenStudy (anonymous):

3. (1 pt) Which number can replace the variable to make these fractions equivalent? n/9 = 56/63 A. 2 B. 7 C. 8 D. 9

OpenStudy (anonymous):

4. (1 pt) Which is a correct comparison of the fractions? 7/12 and 11/20 A. 7/12 < 11/20 B. 7/12 > 11/20 C. 7/12 = 11/20

OpenStudy (anonymous):

5. (1 pt) Which pair of fractions shows 3/4 and 5/6 correctly converted with their least common denominator? A. 3/24 and 5/24 B. 18/24 and 20/24 C. 9/12 and 10/12 D. 27/36 and 30/36

OpenStudy (anonymous):

@YoungStudier help please

OpenStudy (abbycross167):

1 is D

OpenStudy (anonymous):

ok

OpenStudy (abbycross167):

2 is A

OpenStudy (anonymous):

For the first one, you need to see how many times you have to multiply both 7 and 10, and they finally meet up at 70, because 7x10 is 70 and 10x7 is 70.

OpenStudy (abbycross167):

3 is c

OpenStudy (anonymous):

thanks

OpenStudy (abbycross167):

4 is b

OpenStudy (anonymous):

thank you

OpenStudy (abbycross167):

and 5 is c

OpenStudy (abbycross167):

you are very welcome

OpenStudy (anonymous):

ok thank do you want to tell you my score

OpenStudy (anonymous):

AWWWWWWWW 100%

OpenStudy (abbycross167):

Awesome!!!! way to go!